two circular gears whose teeth interlock when the gears turn. Gear A has 72 teeth and gear B has 48 teeth. How many complete revolutions does gear A make when gear B makes 9 complete revolutions?
Answer by mananth(16946)

Gear A 72
gear B = 48
..
Ratio of teeth = 72/48 = 3/2
gear B = 9 rev
..
3 rev of B = 2 rev of A
9 rev of B
9*2/3 = 18 /3
= 6 revolutions